OpenStudy (anonymous):

Use the sigma notation to write this summation in compact form: 1+4+9+16 how would i do this? i see the pattern is +3,+5,+7, etc..... :/ thank you! :)

OpenStudy (anonymous):

would this be right so far? :/ 

hartnn (hartnn):

yeah, now just write the general n'th term formula for the sequence

hartnn (hartnn):

do you see any pattern in the terms ?

OpenStudy (anonymous):

they're adding odd numbers?:/ not quite sure how to interpret the pattern into an equation though:(

hartnn (hartnn):

what about squares ? :P 1^2, 2^2,3^2,...

OpenStudy (anonymous):

ohh i see.. so it's going 1^2 = 1 2^2 = 4 3^2 = 9 4^2 = 16 ? :O so would it be n^2 ? :/ or is there another way to write it?

hartnn (hartnn):

n^2 is totally correct :)

hartnn (hartnn):



OpenStudy (anonymous):

ahh okay yay!! so with sigma notation, you always use the variable k?

hartnn (hartnn):

your wish, you can use 'n' or 'k' or even long names like "\(i\heartsuit food\)"

OpenStudy (anonymous):

hahaha okay:) thank you!! :D i have two more that are for writing summations into compact form... could i post here? or would you be able to help in a new post?

hartnn (hartnn):

new post :)

OpenStudy (anonymous):

okay:)